OpenStudy (anonymous):

4d+9d-(7-6d)=4d+9d-7-6d...how do i solve???

OpenStudy (anonymous):

cancel what is d sym on both side n u gonna b left with 6d=6d

OpenStudy (anonymous):

actually it is wrong 6d is not = to -6d.meaning the question is undefined

OpenStudy (anonymous):

my medal plz

OpenStudy (radar):

Hey you're right, d=0 19d-7 =7d-7 12d=0 d=0 It is not an identity.